[Hallc_running] [New Logentry] RC report
arunts at jlab.org
arunts at jlab.org
Tue Nov 16 09:15:01 EST 2021
After RF work early during the day shift, we resumed data taking at around 11:00 am. Dave Mack performed BCM calibrations at 2:30 pm and we resumed data taking at ~4:00 pm. Later during the day, we had recurring problems with the readout of the HMS Diplole NMR as it just wouldn't lock. After calling Steve Lassiter, and having him walk us through the instructions on switching probes (from probe 5 to 4 and then back to 5), we experienced the same read back error; the probe scans the entire limit range of the probe and and then circles back. Steve Lassiter suggested that as long as the dipole current is stable, we could continue collecting production data (https://logbooks.jlab.org/entry/3944101). The alarms for the readout were disabled and shifters were asked to monitor the dipole current (variable name: ecDI_I_True) on MyaPlot (https://logbooks.jlab.org/entry/3944135). When we change configurations later today, Steve Lassiter will remotely power cycle the NMR and hopefully t!
his will fix the readout problem.
At 11:00 AM, beam will be taken away from us for 15 mins for the injector group would like to investigate whether or not the 0R01 BLM issues originate in the injector. Other that that, we continue to sail through our run plan (currently on page 22/27)!
Let's have an RC meeting today at 10:00 AM EST.
Connection information: https://bluejeans.com/861439920
Highlights and useful links:
How to switch between NMR probes: https://logbooks.jlab.org/entry/3944101
HMS NMR Dipole locking problems -- ALARM DISABLED IN THE ALARM HANDLER: https://logbooks.jlab.org/entry/3944135
This is a plain text email for clients that cannot display HTML. The full logentry can be found online at https://logbooks.jlab.org/entry/3944406
-------------- next part --------------
An HTML attachment was scrubbed...
More information about the Hallc_running